OpenStudy (anonymous):

A car is purchased for $18,000 . After each year, the resale value decreases by 35% . What will the resale value be after 3 years?

OpenStudy (adi3):

first u should multiply 35% * 3

OpenStudy (adi3):

because it is decreasing each year.

OpenStudy (coconutjj):

100% - 35% = 65% A = 18000(0.65)^t Where t is the time.

OpenStudy (coconutjj):

I think this should be correct... I'm not really good with logarithmic equations

OpenStudy (adi3):

but i think fist u should * = 35 * 3 = 105 and then divide 18000/105 = 171.42

OpenStudy (anonymous):

Thank you so much everyone!